2.43x + 13.71 = 8.12x - 22.54
Answer by checkley75(3666)

2.43X+13.71=8.12X-22.54
2.43X-8.12X=-22.54-13.71
-5.69X=-36.25
X=-36.25/-5.69
X=6.37